Price

Quantity demanded

Elasticity coefficient

10

10,000

-

9

13,000

-2.47 

8

17,000


7

22,000


6

25,000



use the arc or midpoint elasticity formula to calculate the elasticity coefficients when price changes from

  1. 10 to 9
  2. 9 to 8
  3. 8 to 7
  4. 7 to 6







1
Expert's answer
2022-02-23T12:23:18-0500

Midpoint elasticity:

"=\frac{\frac{Q_2-Q_1}{Q_2+Q_1}÷2}{\frac{P_2-P_1}{P_2+P_1}÷2}"

a)

"=\frac{\frac{13000-10000}{13000+100000}÷2}{\frac{9-10}{9+10}÷2}=-2.47"

b)

"=\frac{\frac{17000-13000}{17000+13000}÷2}{\frac{8-9}{8+9}÷2}=-2.27"

c)

"=\frac{\frac{22000-17000}{22000+17000}÷2}{\frac{7-8}{8+7}÷2}=-1.92"

d)

"=\frac{\frac{25000-22000}{25000+22000}÷2}{\frac{6-7}{6+7}÷2}=-0.192"